OpenStudy (anonymous):

F(x)= x^2-4x+5 F(1)= F(-1)= F(2)=

OpenStudy (agent47):

just plug in the values in the parenthesis instead of x, and evaluate.

OpenStudy (anonymous):

Thats what i tried doing. For f(1) i got 2 for f(-1) i got 8, but 8 isnt any of my answer choices.

OpenStudy (agent47):

First one: (1)^2-4*(1)+5 and just do the same thing for everything else.

OpenStudy (agent47):

(-1)^2-(4)(-1)+5=1+4+5=10

OpenStudy (anonymous):

You got f(1) correct. For f(-1), you probably made a mistake squaring -1 (-1)^2 = (-1)(-1) = 1

OpenStudy (anonymous):

So for f(2) is that 1?

OpenStudy (anonymous):

Yes =) Good work.

OpenStudy (anonymous):

;) so what about f(a) and f(a+1) ?

OpenStudy (anonymous):

Plug in a and see what you get. Or plug in (a+1) and see what you get. You won't be able to get a single number like before, which is okay.

OpenStudy (anonymous):

So, a^2-4a+5 ?

OpenStudy (anonymous):

Good. Now, when you plug in (a+1) remember to put parenthesis around it.

OpenStudy (anonymous):

do i add the 1 to anything from (a+1) ?

OpenStudy (anonymous):

Plug in (a+1), and then simplify. Let me see your "plug it in" step.

OpenStudy (anonymous):

(a+1)^2-4(a+1)+5 ?

OpenStudy (anonymous):

Perfect. Okay, now you just have to be careful that you simplify correctly =) There are two little tricks to doing that. The first is to remember to distribute the -4. The second is that (a+1)^2 mean (a+1) times itself, so (a+1)(a+1)

OpenStudy (anonymous):

Let me see you try to simplify.

OpenStudy (anonymous):

Um, a^2-2a+15?

OpenStudy (anonymous):

hmmm I'd need to see your middle steps. The 15 is kind of baffling me.

OpenStudy (theviper):

Just put the value x in each case & solve ur prob:)